verify identity.
cosx+1= sinx
Answer #1

cosa al = sina 20 Boost Xlz - ksinacos/ Cosa - Sinar] 7 / 2 = 0 (or cos 1 / 2 = sin I 2 / 2 = 1 1 1 (or) & tant (on 7 / 2 = 1